# Ensemble Kalman Diffusion Guidance: A Derivative-free Method for Inverse Problems

### v1

@@ 4.3 Black-hole imaging inverse problem
The black hole interferometric imaging system aims to reconstruct image of black holes using a set of telescopes distributed on the Earth. Each pair of telescopes produces a measurement $V_{a,b}^{t}$ called visibility, where $(a,b)$ is a pair of telescopes and $t$ is the measuring time. To mitigate the effect of measurement noise caused by atmosphere turbulence and thermal noise, multiple visibilities can be grouped together to cancel out noise (Chael et al., 2018). This produces noise-invariant measurements, termed closure phases ${\bm{y}}^{\mathrm{cph}}_{t,(a,b,c)}$ and log closure amplitudes ${\bm{y}}^{\mathrm{camp}}_{t,(a,b,c,d)}$. We specify the likelihood of this measuring function similar as Sun & Bouman (2021), i.e.,
@@
where $\mathcal{A}_{t}^{\mathrm{cph}}$ and $\mathcal{A}_{t}^{\mathrm{camp}}$ measures the closure phase and log closure amplitude of black hole images ${\bm{x}}$. $\beta_{\mathrm{cph}}$ and $\beta_{\mathrm{camp}}$ are some known parameters given by the telescope system. The first two terms are the sum of chi-square values for closure phases and log closure amplitudes, and the last term constrain the total flux of the black-hole image. We trained a diffusion model on GRMHD dataset (Wong et al., 2022) with resolution 64$\times$ 64 to generate black hole images from telescope measurements.
@@ Results
Figure 4 shows the reconstructed images of the black-hole using our method and the baseline methods with black box access. EnKG is able to generate black hole images with visual features consistent with the ground truth. Moreover, we report quantitative results of each method in Table 3. EnKG achieves relatively low measurement error and the best (blurred) PSNR compared with baseline methods. SCG achieves slightly better data fitting metrics measured by chi-square value, but produces images that are much noisier than those of EnKG.
@@
5 Limitations and discussion
The first limitation of the proposed EnKG is that as a optimization-based method, it cannot capture the exact posterior distribution and thus cannot provide reliable uncertainty quantification, which might be important in some applications. A potential future improvement can involve adapting ideas from the sequential Monte Carlo approach to EnKG to enable provable posterior sampling. Secondly, while the per-sample time cost of EnKG is smaller than the standard gradient-based approach, the total runtime is much longer because EnKG maintains a whole ensemble of interacting particles at each iteration. However, as we shown in Figure 6, even a small number of particles can achieve 20-30% relative $L^{2}$ error. A potential improvement could be introducing a scheme that adaptively controls the number of particles according to the optimization landscape to improve efficiency.
6 Conclusion
@@
Acknowledgments
We thank Ben Prather, Abhishek Joshi, Vedant Dhruv, C.K. Chan, and Charles Gammie for the synthetic images used here, generated under NSF grant AST 20-34306. This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. This research used resources of the Argonne Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC02-06CH11357. This research was done using services provided by the OSG Consortium, which is supported by the National Science Foundation awards #2030508 and #1836650. This research is part of the Delta research computing project, which is supported by the National Science Foundation (award OCI 2005572), and the State of Illinois. Delta is a joint effort of the University of Illinois at Urbana-Champaign and its National Center for Supercomputing Applications.
@@ A.2 Proofs
Under Assumption 12 and 3, suppose the algorithm is implemented with $w_{i}=1/\left(tr\left(C_{yy}^{(i)}\right)\right)$, then
$tr\left(C_{xx}^{(i+1)}\right)\leq tr\left((C_{xx}^{(i)}\right)-\frac{1}{M_{3}}tr\left(C_{xy}^{(i)}C_{xy}^{(i)\top}\right),$ (19)
where
$C_{xy}^{(i)}=\frac{1}{J}\sum_{j=1}^{J}\left({\bm{x}}_{i}^{(j)}-\bar{{\bm{x}}}_{i}\right)\left(\psi({\bm{x}}_{i}^{(j)})-\bar{\psi}_{i}\right)^{\top}.$
Note that $tr\left(C_{xy}^{(i)}C_{xy}^{(i)\top}\right)$ is zero if and only if all ensemble members collapse to a single point, i.e., $tr\left(C_{xx}^{(i)}\right)=0$. Therefore, $tr\left(C_{xx}^{(i)}\right)$ monotonically decreases to zero in the limit as $i$ goes to infinity.
@@
We first rewrite the ensemble update at iteration $i$ as follows
${\bm{x}}_{i+1}^{(j)}={\bm{x}}_{i}^{(j)}+w_{i}C_{xy}^{(i)}\left(y-\psi({\bm{x}}_{i}^{(j)})\right),$ (20)
@@
$C_{xx}^{(i+1)}$ $=\frac{1}{J}\sum_{j}\left[({\bm{x}}_{i}^{(j)}-\bar{{\bm{x}}}_{i})+w_{i}C_{xy}^{(i)}\left(\bar{\psi}_{i}-\psi({\bm{x}}_{i}^{(j)})\right)\right]\left[({\bm{x}}_{i}^{(j)}-\bar{{\bm{x}}}_{i})+w_{i}C_{xy}^{(i)}(\bar{\psi}_{i}-\psi({\bm{x}}_{i}^{(j)}))\right]^{\top}$
$=C_{xx}^{(i)}-2w_{i}C_{xy}^{(i)}C_{xy}^{(i)\top}+w_{i}^{2}C_{xy}^{(i)}C_{yy}^{(i)}C_{xy}^{(i)\top}.$
By linearity of trace, we have
@@
$=tr\left(C_{xx}^{(i)}\right)-\frac{1}{tr\left(C_{yy}^{(i)}\right)}tr\left(C_{xy}^{(i)}C_{xy}^{(i)\top}\right).$
By Assumption 1 and 2, we know there exists a constant $M_{4}>0$ such that $tr\left(C_{yy}^{(i)}\right)\leq M_{4}$. Therefore,
$tr\left(C_{xx}^{(i+1)}\right)\leq tr\left(C_{xx}^{(i)}\right)-\frac{1}{tr\left(C_{yy}^{(i)}\right)}tr\left(C_{xy}^{(i)}C_{xy}^{(i)\top}\right)\leq tr\left(C_{xx}^{(i)}\right)-\frac{1}{M_{4}}tr\left(C_{xy}^{(i)}C_{xy}^{(i)\top}\right).$
By Assumption 3, $tr\left(C_{xy}^{(i)}C_{xy}^{(i)\top}\right)$ is zero if and only if all the ensemble members collapse to a single point, which is $tr\left(C_{xx}^{(i)}\right)=0$. Therefore, $tr\left(C_{xx}^{(i)}\right)$ monotonically decreases to zero. Additionally, we empirically check how quickly the average distance decays as we iterate in our experiments as shown in Figure 5. ∎
Figure 5: Distance of ensemble members quickly decays over update steps. Empirical verification of Lemma 1.
@@
Under Assumption 12 and 3, suppose the algorithm is implemented with $w_{i}<1/\left(tr\left(C_{yy}^{(i)}\right)\right)$, then we have the following approximation:
$C_{xx}^{(i)}\nabla\log\hat{p}\left({\bm{y}}|{\bm{x}}_{i}^{\prime(j)}\right)\approx\frac{1}{J}\sum_{k=1}^{J}\left\langle G(\hat{{\bm{x}}}^{\prime(k)}_{N})-\bar{G},{\bm{y}}-G(\hat{{\bm{x}}}^{(j)}_{N})\right\rangle_{\Gamma}\left({\bm{x}}^{(j)}_{i}-\bar{{\bm{x}}_{i}}\right),$ (22)
@@
$\bar{G}=\frac{1}{J}\sum_{j=1}^{J}G(\hat{{\bm{x}}}^{(j)}_{N})$
@@
$=-\frac{1}{J^{2}}\sum_{k=1}^{J}\sum_{l=1}^{J}\left({\bm{x}}_{i}^{\prime(j)}-\bar{{\bm{x}}}_{i}\right)\left({\bm{D}}\psi\left({\bm{x}}_{i}^{\prime(j)}\right)\left({\bm{x}}_{i}^{\prime(k)}-{\bm{x}}_{i}^{(l)}\right)\right)^{\top}\left(\psi\left({\bm{x}}_{i}^{\prime(j)}\right)-{\bm{y}}\right).$ (27)
@@
which represents the average distance between ensemble members. By Lemma 1, we know that $tr\left(C_{xx}^{(i)}\right)$ monotonically decreases to zero in the limit. Therefore, the ensemble members will get sufficiently close as we iterate. Therefore, we can apply first-order Taylor approximation to $h$ at ${\bm{x}}_{i}^{\prime(j)}$ under Assumption 1 and obtain
@@
where $k\in\{1,\dots,J\}$. Therefore for any $k,l\in\{1,\dots,J\}$, we have
$\psi\left({\bm{x}}_{i}^{\prime(k)}\right)-\psi\left({\bm{x}}_{i}^{\prime(l)}\right)={\bm{D}}\psi\left({\bm{x}}_{i}^{\prime(j)}\right)\left({\bm{x}}_{i}^{\prime(k)}-{\bm{x}}_{i}^{\prime(l)}\right)$
@@
$\approx-\frac{1}{J^{2}}\sum_{k=1}^{J}\sum_{l=1}^{J}\left({\bm{x}}_{i}^{(j)}-\bar{{\bm{x}}}_{i}\right)\left(\psi\left({\bm{x}}_{i}^{\prime(k)}\right)-\psi\left({\bm{x}}_{i}^{(l)}\right)\right)^{\top}\left(\psi\left({\bm{x}}_{i}^{\prime(j)}\right)-{\bm{y}}\right)$
$=-\frac{1}{J}\sum_{k=1}^{J}\left({\bm{x}}_{i}^{\prime(j)}-\bar{{\bm{x}}}_{i}\right)\left(\psi\left({\bm{x}}_{i}^{\prime(k)}\right)-\bar{\psi}_{i}\right)^{\top}\left(\psi\left({\bm{x}}_{i}^{\prime(j)}\right)-{\bm{y}}\right)$
$=-\frac{1}{J}\sum_{k=1}^{J}\left\langle\psi\left({\bm{x}}_{i}^{\prime(k)}\right)-\bar{\psi}_{i},\psi\left({\bm{x}}_{i}^{\prime(j)}\right)-{\bm{y}}\right\rangle\left({\bm{x}}_{i}^{\prime(j)}-\bar{{\bm{x}}}_{i}\right)$
$=\frac{1}{J}\sum_{k=1}^{J}\left\langle G(\hat{{\bm{x}}}^{\prime(k)}_{N})-\bar{G},{\bm{y}}-G(\hat{{\bm{x}}}^{(j)}_{N})\right\rangle\left({\bm{x}}^{(j)}_{i}-\bar{{\bm{x}}_{i}}\right),$

### v2

@@ 5.3 Black-hole imaging inverse problem
The black hole interferometric imaging system aims to reconstruct image of black holes using a set of telescopes distributed on the Earth. Each pair of telescopes produces a measurement $V_{a,b}^{t}$ called visibility, where $(a,b)$ is a pair of telescopes and $t$ is the measuring time. To mitigate the effect of measurement noise caused by atmosphere turbulence and thermal noise, multiple visibilities can be grouped together to cancel out noise (Chael et al., 2018), producing noise-invariant measurements, termed closure phases ${\bm{y}}^{\mathrm{cph}}_{t,(a,b,c)}$ and log closure amplitudes ${\bm{y}}^{\mathrm{camp}}_{t,(a,b,c,d)}$. We specify the likelihood of these measurements similar to Sun & Bouman (2021):
@@
where $\mathcal{A}_{t}^{\mathrm{cph}}$ and $\mathcal{A}_{t}^{\mathrm{camp}}$ measures the closure phase and log closure amplitude of black hole images ${\bm{x}}$. $\beta_{\mathrm{cph}}$ and $\beta_{\mathrm{camp}}$ are known parameters from the telescope system. The first two terms are the sum of chi-square values for closure phases and log closure amplitudes, and the last term constrains the total flux of the black-hole image. We trained a diffusion model on the GRMHD dataset (Wong et al., 2022) with resolution 64$\times$ 64 to generate black hole images from telescope measurements.
@@ Results
Figure 4 shows the reconstructed images of the black-hole using our EnKG method and the baseline methods with black box access. EnKG is able to generate black hole images with visual features consistent with the ground truth. Table 3 shows the quantitative comparison. EnKG achieves relatively low measurement error (i.e., consistency with observations) and the best (blurred) PSNR compared with baseline methods (i.e., realistic images). SCG achieves slightly better data fitting metrics, but produces much noisier images than those by EnKG (Figure 4).
@@
6 Conclusion and discussion
@@
Despite its strengths, EnKG has certain limitations. First, as an optimization-based approach, it does not aim to recover the full posterior distribution and therefore cannot provide reliable uncertainty quantification—an important feature in some applications. Second, while EnKG reduces per-sample computational cost compared to standard gradient-based methods, its total runtime is higher due to maintaining and updating an ensemble of interacting particles. Future work could explore adaptive strategies to dynamically adjust the number of particles based on the optimization landscape to improve efficiency. Additionally, integrating fast diffusion model sampling techniques (Zheng et al., 2023; Song et al., 2023c; Yin et al., 2024) may further reduce computational overhead.
@@ A.1 Notation
$\Gamma$ Covariance matrix of the Gaussian noise model
$\left\langle\cdot,\cdot\right\rangle_{\Gamma}$ Weighted Euclidean inner product, $\left\langle\cdot,\cdot\right\rangle_{\Gamma}=\left\langle\cdot,\Gamma^{-1}\cdot\right\rangle$
@@ A.2 Proofs
Under Assumption 12 and 3, suppose the correction step is implemented with $w_{i}=1/\left(tr\left(C_{yy}^{(i)}\right)\right)$ as follows,
${\bm{x}}_{i+1}^{(j)}={\bm{x}}_{i}^{(j)}+w_{i}C_{xy}^{(i)}\left({\bm{y}}-\psi\left({\bm{x}}_{i}^{(j)}\right)\right),$ (23)
where $j\in\{1,\dots,J\}$ and
$C_{xy}^{(i)}=\frac{1}{J}\sum_{j=1}^{J}\left({\bm{x}}_{i}^{(j)}-\bar{{\bm{x}}}_{i}\right)\left(\psi\left({\bm{x}}_{i}^{(j)}\right)-\bar{\psi}_{i}\right)^{\top}.$
Then $tr\left(C_{xx}^{(i)}\right)$ monotonically decreases to zero in the limit as $i$ goes to infinity.
@@
We first start from the ensemble update of the correction step given in Eq. (23) at iteration $i$ as follows
${\bm{x}}_{i+1}^{(j)}={\bm{x}}_{i}^{(j)}+w_{i}C_{xy}^{(i)}\left({\bm{y}}-\psi\left({\bm{x}}_{i}^{(j)}\right)\right),$ (24)
@@
$C_{xx}^{(i+1)}=$ $\frac{1}{J}\sum_{j=1}^{J}\left[({\bm{x}}_{i}^{(j)}-\bar{{\bm{x}}}_{i})+w_{i}C_{xy}^{(i)}\left(\bar{\psi}_{i}-\psi({\bm{x}}_{i}^{(j)})\right)\right]\left[({\bm{x}}_{i}^{(j)}-\bar{{\bm{x}}}_{i})+w_{i}C_{xy}^{(i)}(\bar{\psi}_{i}-\psi({\bm{x}}_{i}^{(j)}))\right]^{\top}$
$=$ $\frac{1}{J}\sum_{j=1}^{J}\left[({\bm{x}}_{i}^{(j)}-\bar{{\bm{x}}}_{i})({\bm{x}}_{i}^{(j)}-\bar{{\bm{x}}}_{i})^{\top}+w_{i}^{2}C_{xy}^{(i)}\left(\bar{\psi}_{i}-\psi({\bm{x}}_{i}^{(j)})\right)\left(\bar{\psi}_{i}-\psi({\bm{x}}_{i}^{(j)})\right)^{\top}C_{xy}^{(i)\top}\right]$
$+\frac{1}{J}\sum_{j=1}^{J}\left[w_{i}C_{xy}^{(i)}\left(\bar{\psi}_{i}-\psi({\bm{x}}_{i}^{(j)})\right)({\bm{x}}_{i}^{(j)}-\bar{{\bm{x}}}_{i})^{\top}+w_{i}({\bm{x}}_{i}^{(j)}-\bar{{\bm{x}}}_{i})(\bar{\psi}_{i}-\psi({\bm{x}}_{i}^{(j)}))^{\top}C_{xy}^{(i)\top}\right].$ (26)
We notice that
$\frac{1}{J}\sum_{j=1}^{J}w_{i}C_{xy}^{(i)}\left(\bar{\psi}_{i}-\psi({\bm{x}}_{i}^{(j)})\right)\left({\bm{x}}_{i}^{(j)}-\bar{{\bm{x}}}_{i}\right)^{\top}$ $=-w_{i}C_{xy}^{(i)}C_{xy}^{(i)\top}$
$\frac{1}{J}\sum_{j=1}^{J}w_{i}\left({\bm{x}}_{i}^{(j)}-\bar{{\bm{x}}}_{i}\right)\left(\bar{\psi}_{i}-\psi({\bm{x}}_{i}^{(j)})\right)^{\top}C_{xy}^{(i)\top}$ $=-w_{i}C_{xy}^{(i)}C_{xy}^{(i)\top}.$
Therefore, we can rewrite Eq. (A.2) as follows:
$C_{xx}^{(i+1)}=C_{xx}^{(i)}-2w_{i}C_{xy}^{(i)}C_{xy}^{(i)\top}+w_{i}^{2}C_{xy}^{(i)}C_{yy}^{(i)}C_{xy}^{(i)\top}.$
Further, by linearity of trace, we have
@@
$=tr\left(C_{xx}^{(i)}\right)-\frac{1}{tr\left(C_{yy}^{(i)}\right)}tr\left(C_{xy}^{(i)}C_{xy}^{(i)\top}\right).$ (27)
By Assumption 1 and 2, we know that both $tr\left(C_{xx}^{(i)}\right)$ and $tr\left(C_{yy}^{(i)}\right)$ are upper bounded. By Assumption 3, $tr\left(C_{xy}^{(i)}C_{xy}^{(i)\top}\right)$ is lower bounded unless all the ensemble members collapse to a single point. Thus, there exists a $\alpha>0$ such that $tr\left(C_{xy}^{(i)}C_{xy}^{(i)\top}\right)\geq\alpha\cdot tr\left(C_{xx}^{(i)}\right)tr\left(C_{yy}^{(i)}\right)$. Therefore,
$tr\left(C_{xx}^{(i+1)}\right)\leq tr\left(C_{xx}^{(i)}\right)-\frac{1}{tr\left(C_{yy}^{(i)}\right)}tr\left(C_{xy}^{(i)}C_{xy}^{(i)\top}\right)\leq\left(1-\alpha\right)tr\left(C_{xx}^{(i)}\right).$
Note that from Eq. (27), we have $\alpha\leq 1$. Therefore, $tr\left(C_{xx}^{(i)}\right)$ monotonically decreases to zero. Additionally, we empirically check how quickly the average distance decays as we iterate in our experiments as shown in Figure 5. ∎
Figure 5: Empirical verification of Lemma 1 and Assumption 3. Top: the distance of ensemble members quickly decays over update steps. Bottom: while larger $tr(C_{xx})$ does not always indicate larger $tr(C_{yy})$, $tr(C_{yy})$ is close zero only when $tr(C_{xx})$ is close to zero.
@@
Under Assumption 12 and 3, suppose the correction step is implemented as follows with $w_{i}=1/\left(tr\left(C_{yy}^{(i)}\right)\right)$,
${\bm{x}}_{i+1}^{(j)}$ $={\bm{x}}_{i}^{\prime(j)}+w_{i}C_{xy}^{(i)}\left({\bm{y}}-\psi\left({\bm{x}}_{i}^{\prime(j)}\right)\right)$ (28)
$={\bm{x}}_{i}^{\prime(j)}+w_{i}\frac{1}{J}\sum_{k=1}^{J}\left\langle\psi\left({\bm{x}}^{\prime(k)}_{i}\right)-\bar{G},{\bm{y}}-\psi\left({\bm{x}}^{\prime(j)}_{i}\right)\right\rangle_{\Gamma}\left({\bm{x}}^{\prime(j)}_{i}-\bar{{\bm{x}}_{i}}\right),$ (29)
@@
$C_{xy}^{(i)}=\frac{1}{J}\sum_{j=1}^{J}\left({\bm{x}}_{i}^{\prime(j)}-\bar{{\bm{x}}}_{i}\right)\left(\psi\left({\bm{x}}_{i}^{\prime(j)}\right)-\bar{\psi}_{i}\right)^{\top}.$
After sufficient iterations, we have the following approximation:
$C_{xy}^{(i)}\left({\bm{y}}-\psi\left({\bm{x}}_{i}^{\prime(j)}\right)\right)$ $=\frac{1}{J}\sum_{k=1}^{J}\left\langle\psi\left({\bm{x}}^{\prime(k)}_{i}\right)-\bar{G},{\bm{y}}-\psi\left({\bm{x}}^{\prime(j)}_{i}\right)\right\rangle_{\Gamma}\left({\bm{x}}^{\prime(j)}_{i}-\bar{{\bm{x}}_{i}}\right)$ (30)
$\approx C_{xx}^{(i)}\nabla_{{\bm{x}}}\log\hat{p}\left({\bm{y}}|{\bm{x}}_{i}^{\prime(j)}\right).$ (31)
@@
$=-\frac{1}{J^{2}}\sum_{k=1}^{J}\sum_{l=1}^{J}\left({\bm{x}}_{i}^{\prime(k)}-\bar{{\bm{x}}}_{i}\right)\left({\bm{D}}\psi\left({\bm{x}}_{i}^{\prime(j)}\right)\left({\bm{x}}_{i}^{\prime(k)}-{\bm{x}}_{i}^{(l)}\right)\right)^{\top}\left(\psi\left({\bm{x}}_{i}^{\prime(j)}\right)-{\bm{y}}\right).$ (36)
@@
which represents the average distance between ensemble members. By Lemma 1, we know that $tr\left(C_{xx}^{(i)}\right)$ monotonically decreases to zero in the limit. Therefore, the ensemble members will get sufficiently close as we iterate. Therefore, we can apply first-order Taylor approximation to $\psi$ at ${\bm{x}}_{i}^{\prime(j)}$ under Assumption 1 and obtain
@@
where $k\in\{1,\dots,J\}$. Therefore for any $k,l\in\{1,\dots,J\}$, by applying the approximation above at both ${\bm{x}}_{i}^{\prime(k)}$ and ${\bm{x}}_{i}^{\prime(l)}$, we have
$\psi\left({\bm{x}}_{i}^{\prime(k)}\right)-\psi\left({\bm{x}}_{i}^{\prime(l)}\right)\approx{\bm{D}}\psi\left({\bm{x}}_{i}^{\prime(j)}\right)\left({\bm{x}}_{i}^{\prime(k)}-{\bm{x}}_{i}^{\prime(l)}\right)$
@@
$\approx-\frac{1}{J^{2}}\sum_{k=1}^{J}\sum_{l=1}^{J}\left({\bm{x}}_{i}^{\prime(k)}-\bar{{\bm{x}}}_{i}\right)\left(\psi\left({\bm{x}}_{i}^{\prime(k)}\right)-\psi\left({\bm{x}}_{i}^{(l)}\right)\right)^{\top}\left(\psi\left({\bm{x}}_{i}^{\prime(j)}\right)-{\bm{y}}\right)$
$=-\frac{1}{J}\sum_{k=1}^{J}\left({\bm{x}}_{i}^{\prime(k)}-\bar{{\bm{x}}}_{i}\right)\left(\psi\left({\bm{x}}_{i}^{\prime(k)}\right)-\bar{\psi}_{i}\right)^{\top}\left(\psi\left({\bm{x}}_{i}^{\prime(j)}\right)-{\bm{y}}\right)$
$=-\frac{1}{J}\sum_{k=1}^{J}\left\langle\psi\left({\bm{x}}_{i}^{\prime(k)}\right)-\bar{\psi}_{i},\psi\left({\bm{x}}_{i}^{\prime(j)}\right)-{\bm{y}}\right\rangle\left({\bm{x}}_{i}^{\prime(k)}-\bar{{\bm{x}}}_{i}\right)$
$=\frac{1}{J}\sum_{k=1}^{J}\left\langle G(\hat{{\bm{x}}}^{\prime(k)}_{N})-\bar{G},{\bm{y}}-G(\hat{{\bm{x}}}^{(k)}_{N})\right\rangle\left({\bm{x}}^{\prime(k)}_{i}-\bar{{\bm{x}}_{i}}\right),$